# ChipletPart: Cost-Aware Partitioning for 2.5D Systems

### v2

@@ 6.1. Benchmarks and Baselines
Metrics. For simplicity, we assume iso-performance for our testcases, even across technology nodes; instead, we derive advantages in power from improvements in IO count or changes in technology node. We also assume that connections between blocks in a chiplet will have the same speed as inter-chiplet connections, although the inter-chiplet connections will have higher power and area requirements due to large IO drivers. Design area is a major contributor to cost, and chip power also contributes to chip cost via added bumps (for current delivery) that require additional area. Given this, we use cost as a single metric for partitioning.
Baselines. Since there is no open-source, cost-driven chiplet partitioner available, we compare against: (i) Min-cut partitioners (hMETIS (KarypisK96, ), which is often used in chiplet flows (Kabir20, ), and TritonPart (BustanyGKKPW23, )) applied to weighted hypergraphs, where each hyperedge is weighted by the corresponding I/O bandwidth; (ii) the parChiplet method from FloorPlet (ChenLZZL23, ); and (iii) partitioning solutions generated by human engineers.161616 Our attempts to contact the authors of (VicharraCH25, ) were unsuccessful. We are currently in correspondence with the authors of (Li24, ), but have encountered difficulties running their implementation. Efforts to resolve these issues are ongoing.
Table 5. Partitioning Results for hMETIS, TritonPart, FloorPlet (ChenLZZL23, ), Manual, and ChipletPart. ChipletPart No-FP has no floorplan checker. “Feas” indicates IO-feasibility (✓: feasible, X: not feasible). Best (feasible) cost is in bold.
Benchmark hMETIS TritonPart FloorPlet Manual ChipletPart No-FP ChipletPart $|\mathcal{C}|$ Cost Feas $|\mathcal{C}|$ Cost Feas $|\mathcal{C}|$ Cost Feas $|\mathcal{C}|$ Cost Feas $|\mathcal{C}|$ Cost Feas $|\mathcal{C}|$ Cost Feas WS1 5 72.4 X 3 77.3 6 55.5 X 1 71.2 7 55.1 X 6 53.9 WS2 8 138.1 X 7 143.2 X 4 150.3 2 145.5 6 122.3 X 8 123.4 WS3 7 316.3 X 8 362.4 X 4 310.5 4 310.5 8 319.6 X 7 300.4 WS4 8 1449.5 X 8 1564.2 X 5 1237.9 X 8 674.3 7 776.2 X 8 659.9 MP 1 5.7 1 5.7 1 5.7 1 5.7 1 5.7 1 5.7 TC1 3 72.6 3 68.1 6 45.6 X 1 70.5 4 47.2 6 46.5 TC2 3 73.9 3 71.1 7 46.5 X 4 49.4 6 48.7 X 7 46.2 EPYC 4 92.1 4 96.3 4 89.3 4 80.8 8 76.1 X 8 72.5 GA100 3 52.3 3 60.5 2 42.1 5 36.1 2 33.6 5 31.8
6.2. Validations of Chiplet Partitioning
Given that previous works use traditional min-cut partitioners in 2.5D flows, we directly compare ChipletPart with leading partitioners hMETIS and TritonPart. We also compare to the Floorplet floorplan-aware chiplet partitioner (ChenLZZL23, ) and manual partitions. For the waferscale testcases based on (Pal2021, ) we compare the memory-compute partition in that work with a per-tile partitioning as the manual partition. Overall, our validations span (i) validation of chiplet cost-driven partitioning, (ii) validation of floorplan-awareness, (iii) validation of heterogeneous technology-awareness, and (iv) analyses of hyperparameter sensitivities and runtime.
6.2.1. Validation of chiplet cost-driven partitioning
We evaluate ChipletPart against the baselines listed in the previous section: hMETIS, TritonPart, FloorPlet and manually derived partitions. Our validation is thus divided into three parts, detailed below.
Comparison with netlist partitioners. We evaluate the chiplet cost-driven partitioning capabilities of ChipletPart by benchmarking against leading netlist partitioners hMETIS and TritonPart, using the latter’s default parameter settings. [^17] For hMETIS, the default settings are Nruns = 10, CType = 1, RType = 1, Vcycle = 1, Reconst = 0, and seed = 0 (KarypisAKS99, ). For TritonPart, the default parameter settings are thr_coarsen_hyperedge_size_skip=200thr\_coarsen\_hyperedge\_size\_skip=200, coarsening_ratio=1.6coarsening\_ratio=1.6, max_moves=60max\_moves=60, and num_coarsen_solutions=3num\_coarsen\_solutions=3 (BustanyGKKPW23, ). We use the testcases listed in Table 4 for our evaluations, with results presented in Table 5. For both hMETIS and TritonPart:
  • We set an imbalance factor of 5%, and define our input set of partitions as K=1,2,3,,10{K}={1,2,3,...,10}.
  • We evaluate each partitioning solution using our cost model (Secs. 4 and 4.2), and report the solution with lowest chiplet cost.
  • Since hMETIS and TritonPart are not technology-aware, for a fair comparison, we enforce homogeneity by running ChipletPart with only the $7\text{nm}$ technology node. Consequently, the cost model evaluates all partitioning solutions using the $7\text{nm}$ technology node. The floorplanner is run in the standard mode (Section 5.2).
Results are presented in Table 5. Columns $|C|$, CostCost, and $Feas$ respectively indicate the number of chiplets, the evaluated cost of the chiplet solution, and whether the solution satisfies the reach constraints. ChipletPart consistently produces solutions with better cost than hMETIS and TritonPart, with improvements of up to 55% and 58%, respectively. The geometric mean improvements over hMETIS and TritonPart are 16% and 20%, respectively. This result shows the advantages of ChipletPart’s cost-driven partitioning over traditional min-cut partitioners. Furthermore, ChipletPart consistently produces floorplan-feasible solutions, whereas standard netlist partitioners often fail to meet the reach constraints. 191919hMETIS and TritonPart optimize cutsize and return partitions with better cutsize compared to ChipletPart (albeit with worse chiplet cost).
Comparison with previous work. We compare ChipletPart with the parChiplet partitioner from FloorPlet (ChenLZZL23, ). Here, we set the FloorPlet parameters as $|\mathcal{C}|_{\min}=1$, $|\mathcal{C}|_{\max}=10$ and $rr=1.0$. All partitioning solutions are evaluated using our cost model with the $7\text{nm}$ technology node and the floorplanner is run in standard mode. Table 5 presents these results. Compared to FloorPlet, ChipletPart achieves up to a 47% cost improvement, with a geometric mean improvement of $6\%$. While FloorPlet produces a solution with slightly lower cost on the TC1 testcase, the solution violates reach constraints. In contrast, ChipletPart consistently generates floorplan-feasible solutions with better cost, highlighting its strength in delivering high-quality and physically realizable chiplet partitions.
Comparison with human baselines. We generate human baselines for each testcase as follows. For the Waferscale ($WS_{1},WS_{2},WS_{3},WS_{4}$) designs, which are tile-based, the human partitioning solution aligns directly with these tiles. For the Industry testcases, we split around the crossbar; $TC_{1}$ is a single partition and $TC_{2}$ is a 4-way partition with the same blocks as $TC_{1}$ but with a 4-way split crossbar. For the EPYC testcase (Figure 10), we cluster each set of cores with its corresponding L3 — distributing other blocks to maintain regularity and balanced area — resulting in four partitions each containing 2 PCIe blocks, 1 DDR, 1 L3, and 4 cores. Similarly, for GA100, we create five chiplets: four partitions each holding 32 simultaneous multiprocessors and 10 L2 blocks, and one partition for the remaining blocks. [^20] Our human baselines do not necessarily reflect partitions chosen by industry. Instead, we split the blocks to generate relatively even partitions with small cutset and high regularity as an example of what the human engineer might decide to do.
Similar to the previous sections, all partitioning solutions are evaluated using our cost model with the $7\text{nm}$ technology node and the floorplanner is run in standard mode. Table 5 presents the results. ChipletPart generates solutions that are up to $24\%$ better than the manual baselines with a geometric mean improvement of $7\%$. While the human baselines are all floorplan-feasible, they are significantly suboptimal. We see that a black-box partitioner such as ChipletPart can find better-quality solutions compared to manual strategies.
6.2.2. Validation of floorplan-awareness
We also evaluate the impact of floorplan-awareness in ChipletPart. Table 5 compares solutions generated with and without floorplan-awareness enabled. Disabling floorplan-awareness can result in lower nominal chiplet costs ($WS_{2}$), but frequently produces solutions that violate interconnect reach constraints. In fact, six out of nine testcases produce infeasible solutions when floorplan-awareness is disabled. In addition to feasibility, floorplan-awareness leads to up to $15\%$ improvement in chiplet cost, with a geometric mean improvement of $4\%$. This is because our cost model incorporates floorplan-level information when evaluating partition quality — thus, floorplan-aware solutions better align with the true cost objective. These observations highlight the importance of incorporating floorplan-awareness: while it may slightly increase nominal cost in some cases, it helps to ensure that all generated solutions are physically realizable.
@@ 6.2.3. Validation of heterogeneous technology-awareness
We evaluate benefits of incorporating heterogeneous technology awareness on all nine testcases (Table 4) and a technology list comprising $14\text{nm}$, $10\text{nm}$, and $7\text{nm}$ nodes. Table 6 shows chiplet cost reductions achieved through technology awareness relative to the best solutions obtained via homogeneous technology integration, for each technology node. We first run ChipletPart using a single technology node per testcase (homogeneous integration) to obtain partitioning solutions for each technology. We then run ChipletPart with technology awareness enabled (i.e., leveraging all three technology nodes) and compare the results. Our results show that heterogeneous technology assignment can reduce system cost by up to 43%, with geometric mean cost reductions of 7%, 15%, and 15% when compared against homogeneous solutions at $7\text{nm}$, $10\text{nm}$, and $14\text{nm}$, respectively. [^21] To further validate the effectiveness of our GA, we implemented an enumeration-based framework that exhaustively explores all $\sum_{k=1}^{K_{\text{max}}}\binom{k+m-1}{m-1}$ possible assignments. Our experiments show that the GA achieves up to $7\times$ speedup with less than 1% degradation in chiplet cost compared to the enumeration-based framework. Note, that all solutions generated by ChipletPart are floorplan feasible.
@@
[…]
Figure 12. Sensitivity analysis of hyperparameters. Larger (distance from plot center) indicate greater influence.
[…]
@@ 6.2.4. Hyperparameter exploration and sensitivity analysis
ChipletPart relies on four key hyperparameters for its genetic algorithm (GA): (i) total population size $tot_{pop}$; (ii) maximum number of generations $\Psi$; (iii) mutation probability $p_{m}$; and (iv) crossover probability $p_{c}$ (cf. Algorithm 1). [^22] ChipletPart never returns a chiplet solution with more than 8 chiplets, so we set $K_{max}$ to 8. We validate the default values of these hyperparameters via an empirical study using testcases $WS_{1}$, $WS_{2}$, $TC_{1}$, and $TC_{2}$ from Table 4. We use the following technology nodes for this study: $7\text{nm}$, $10\text{nm}$ and $14\text{nm}$, and run ChipletPart on the four aforementioned testcases. Figure 11 presents the normalized chiplet costs (normalized to the cost of manual solutions) and the normalized runtime (represented as “Normalized Value” in Figure 11), normalized to a constant value of $10000$ seconds.232323 $10000$ seconds is chosen for normalization to allow for comparison with about three hours of runtime. In our experiments, we systematically vary each hyperparameter while keeping the others fixed at their default values. For each configuration, we measure the resulting chiplet cost and runtime across the testcases. Our observations from Figure 11 indicates that the default hyperparameter settings achieve a balanced trade-off between partitioning quality and runtime efficiency.

### v3

@@ 6.1. Benchmarks and Baselines
Metrics. For simplicity, we assume iso-performance for our testcases, even across technology nodes; instead, we derive advantages in power from improvements in I/O count or changes in technology node. We also assume that connections between blocks in a chiplet will have the same speed as inter-chiplet connections, although the inter-chiplet connections will have higher power and area requirements due to large I/O drivers. Design area is a major contributor to cost, and chip power contributes to chip cost via added bumps for current delivery that require additional area. For this reason, we use cost as a single metric for partitioning.
Baselines. Since there is no open-source, cost-driven chiplet partitioner available, we compare against: (i) Min-cut partitioners (hMETIS (KarypisK96, ), which is often used in chiplet flows (Kabir20, ), and TritonPart (BustanyGKKPW23, )) applied to weighted hypergraphs, where each hyperedge is weighted by the corresponding I/O bandwidth; (ii) the parChiplet method from Floorplet (ChenLZZL23, ); (iii) Chipletizer (Li24, ), for which we thank the authors for providing access to their code; and (iv) partitioning solutions generated by human engineers.
Table 5. Partitioning costs and number of chiplets for hMETIS, TritonPart, Floorplet (ChenLZZL23, ), Manual, Chipletizer (Li24, ), $\text{ChipletPart}_{\text{No-FP}}$, and ChipletPart. Best cost per benchmark is in bold. We additionally report the geometric mean (GM) of absolute costs as well as the GM percentage improvement relative to the Manual baseline. Positive GM improvement indicates lower (better) cost on average. Chipletizer results are shown only for testcases where valid solutions were obtained. Because these cover a subset of the full design suite, GM values computed over six testcases are not directly comparable to the nine-testcase GM; hence, we denote Chipletizer’s GM cost (and improvement %) as N/A.1818footnotemark: 18
Benchmark hMETIS TritonPart Floorplet Manual Chipletizer $\text{ChipletPart}_{\text{No-FP}}$ ChipletPart $|\mathcal{C}|$ Cost $|\mathcal{C}|$ Cost $|\mathcal{C}|$ Cost $|\mathcal{C}|$ Cost $|\mathcal{C}|$ Cost $|\mathcal{C}|$ Cost $|\mathcal{C}|$ Cost WS1 5 72.4 3 77.3 6 55.5 1 71.2 1 71.2 7 55.1 6 53.9 WS2 8 138.1 7 143.2 4 150.3 2 145.5 Fail Fail 6 122.3 8 123.4 WS3 7 316.3 8 362.4 4 310.5 4 310.5 Fail Fail 8 319.6 7 300.4 WS4 8 1449.5 8 1564.2 5 1237.9 8 674.3 Fail Fail 7 776.2 8 659.9 MP 1 5.7 1 5.7 1 5.7 1 5.7 1 5.7 1 5.7 1 5.7 TC1 3 72.6 3 68.1 6 45.6 1 70.5 1 70.5 4 47.2 6 46.5 TC2 3 73.9 3 71.1 7 46.5 4 49.4 1 74.9 6 48.7 7 46.2 EPYC 4 92.1 4 96.3 4 89.3 4 80.8 6 139.3 8 76.1 8 72.5 GA100 3 52.3 3 60.5 2 42.1 5 36.1 5 43.8 2 33.6 5 31.8 GM cost 148.3 171.0 113.7 114.9 N/A [^17] For the six testcases where Chipletizer produces solutions, the geometric mean cost (GM cost) is 48.5 vs. 33.9 for ChipletPart, with ChipletPart achieving a geo-mean cost improvement (GM imprv.) of 30.1%. 85.6 82.3 GM impr. (%) -21% -26% -2% 0% N/A [^17] For the six testcases where Chipletizer produces solutions, the geometric mean cost (GM cost) is 48.5 vs. 33.9 for ChipletPart, with ChipletPart achieving a geo-mean cost improvement (GM imprv.) of 30.1%. +9% +13%
Table 6. Chiplet count and I/O-feasibility for hMETIS, TritonPart, Floorplet (ChenLZZL23, ), Manual, Chipletizer (Li24, ), $\text{ChipletPart}_{\text{No-FP}}$, and ChipletPart. “Feas” indicates I/O-feasibility (✓: feasible, X: not feasible; “Fail” indicates that no valid solution was obtained). We also report the number of successful runs for each category.
Benchmark hMETIS TritonPart Floorplet Manual Chipletizer $\text{ChipletPart}_{\text{No-FP}}$ ChipletPart $|\mathcal{C}|$ Feas $|\mathcal{C}|$ Feas $|\mathcal{C}|$ Feas $|\mathcal{C}|$ Feas $|\mathcal{C}|$ Feas $|\mathcal{C}|$ Feas $|\mathcal{C}|$ Feas WS1 5 X 3 6 X 1 1 7 X 6 WS2 8 X 7 X 4 2 Fail Fail 6 X 8 WS3 7 X 8 X 4 4 Fail Fail 8 X 7 WS4 8 X 8 X 5 X 8 Fail Fail 7 X 8 MP 1 1 1 1 1 1 1 TC1 3 3 6 X 1 1 4 6 TC2 3 3 7 X 4 1 6 X 7 EPYC 4 4 4 4 6 X 8 X 8 GA100 3 3 2 5 5 X 2 5 #Successes (out of 9) 4 6 5 9 4 3 9
6.2. Validations of Chiplet Partitioning
Given that previous works use traditional min-cut partitioners in 2.5D flows, we directly compare ChipletPart with leading partitioners hMETIS and TritonPart. We also compare to the Floorplet floorplan-aware chiplet partitioner (ChenLZZL23, ), Chipletizer (Li24, ), and manual partitions. For the Waferscale testcases based on (Pal2021, ) we compare the memory-compute partition in that work with a per-tile partitioning as the manual partition. Overall, our validations span (i) validation of chiplet cost-driven partitioning, (ii) validation of floorplan-awareness, (iii) validation of heterogeneous technology-awareness, and (iv) analyses of hyperparameter sensitivities and runtime.
6.2.1. Validation of chiplet cost-driven partitioning
We evaluate ChipletPart against the baselines listed in the previous section: hMETIS, TritonPart, Floorplet, Chipletizer, and manually derived partitions. Our validation is thus divided into three parts, detailed below.
Comparison with netlist partitioners. We evaluate the chiplet cost-driven partitioning capabilities of ChipletPart by benchmarking against leading netlist partitioners hMETIS and TritonPart, using the latter’s default parameter settings. [^19] For hMETIS, the default settings are Nruns = 10, CType = 1, RType = 1, Vcycle = 1, Reconst = 0, and seed = 0 (KarypisAKS99, ). For TritonPart, the default parameter settings are thr_coarsen_hyperedge_size_skip=200thr\_coarsen\_hyperedge\_size\_skip=200, coarsening_ratio=1.6coarsening\_ratio=1.6, max_moves=60max\_moves=60, and num_coarsen_solutions=3num\_coarsen\_solutions=3 (BustanyGKKPW23, ). We use the testcases listed in Table 4 for our evaluations, with results presented in Tables 18 and 6. Table 18 presents chiplet cost comparisons and Table 6 presents floorplan-feasibility comparisons. For both hMETIS and TritonPart:
  • We set an imbalance factor of 5% and define our input set of partitions as K=1,2,3,,10{K}={1,2,3,...,10}.
  • We evaluate each partitioning solution using our cost model (Secs. 4 and 4.2) and report the solution with lowest chiplet cost.
  • Since hMETIS and TritonPart are not technology-aware, for a fair comparison, we enforce homogeneity by running ChipletPart with only the 7 nm technology node. Consequently, the cost model evaluates all partitioning solutions using the 7 nm technology node. The floorplanner is run in the standard mode (Section 5.2).
Results are presented in Tables 18 and 6. Columns $|C|$, CostCost, and $Feas$, respectively indicate the number of chiplets, the evaluated cost of the chiplet solution, and whether the solution satisfies the reach constraints. ChipletPart consistently produces solutions with better cost than hMETIS and TritonPart with improvements of up to 55% and 58% respectively. The geometric mean improvements over hMETIS and TritonPart are 16% and 20%, respectively. This result shows the advantages of ChipletPart’s cost-driven partitioning over traditional min-cut partitioners. Furthermore, ChipletPart consistently produces floorplan-feasible solutions, whereas standard netlist partitioners often fail to meet the reach constraints.212121hMETIS and TritonPart optimize cutsize and return partitions with better cutsize compared to ChipletPart (albeit with worse chiplet cost).
Comparison with previous work. We compare ChipletPart with the parChiplet partitioner from Floorplet (ChenLZZL23, ), and Chipletizer (Li24, ). Here, we set the Floorplet parameters as $|\mathcal{C}|_{\min}=1$, $|\mathcal{C}|_{\max}=10$ and $rr=1.0$. All partitioning solutions are evaluated using our cost model with the 7 nm technology node and the floorplanner is run in standard mode. Tables 18 and 6 present these results. Compared to Floorplet, ChipletPart achieves up to a 47% cost improvement, with a geometric mean improvement of $6\%$. While Floorplet produces a solution with slightly lower cost on the $TC_{1}$ testcase, the solution violates reach constraints. In contrast, ChipletPart consistently generates floorplan-feasible solutions with better cost, highlighting its strength in delivering high-quality and physically realizable chiplet partitions. For Chipletizer, we use the default settings from the original implementation, with production volume (Li24, ) set to 500,000 units. Chipletizer fails to produce a solution on three of nine testcases ($WS_{1},WS_{2},WS_{3},WS_{4}$) due to scalability limitations. Across the remaining testcases, ChipletPart achieves up to 48% lower cost [^22] A more effective way to run Chipletizer would be to integrate it with our cost model. Since this is beyond the scope of our current work, we leave it for future research. compared to Chipletizer, while consistently generating I/O-feasible solutions. For example, on GA100, both frameworks produce a five-chiplet solution, but ChipletPart achieves a 27% cost reduction and satisfies I/O-feasibility, whereas the Chipletizer solution violates feasibility.
Comparison with human baselines. We generate human baselines for each testcase as follows. For the Waferscale ($WS_{1},WS_{2},WS_{3},WS_{4}$) designs, which are tile-based, the human partitioning solution aligns directly with these tiles. For the red4-way partition with the same blocks as $TC_{1}$ but with a 4-way split crossbar. For the EPYC testcase (Figure 10), we cluster each set of cores with its corresponding L3 — distributing other blocks to maintain regularity and balanced area — resulting in four partitions each containing two PCIe blocks, one DDR, one L3, and four cores. Similarly, for GA100, we create five chiplets: four partitions each holding 32 simultaneous multiprocessors and 10 L2 blocks, and one partition for the remaining blocks. [^23] Our human baselines do not necessarily reflect partitions chosen by industry. Instead, we split the blocks to generate relatively even partitions with small cutset and high regularity as an example of what the human engineer might decide to do.
Similar to the previous sections, all partitioning solutions are evaluated using our cost model with the $7\text{nm}$ technology node and the floorplanner is run in standard mode. Tables 18 and 6 present the results. ChipletPart generates solutions that are up to $34\%$ better than the manual baselines with a geometric mean improvement of $13\%$. While the human baselines are all floorplan-feasible, they are not cost-optimal. These results indicate that a black-box partitioner such as ChipletPart can generate higher-quality solutions than manual partitioning strategies.
6.2.2. Validation of floorplan-awareness
@@
Refer to caption A bar chart showing the performance of running ChipletPart with some of the features turned off.
Figure 11. Impact of initial partitioning and refinement in Core-ChipletPart. Pool-Only denotes Core-ChipletPart with no refinement. KL-Only denotes Core-ChipletPart with KL refinement only. FM-Only denotes Core-ChipletPart with FM refinement only.
6.2.3. Impact of initial partitioning and refinement
To quantify the contribution of each component in Core-ChipletPart, we conduct an ablation study comparing four configurations: (i) the full framework i.e., Core-ChipletPart; (ii) initial partitioning pool-only with no refinement; (iii) initial partitioning pool with FM refinement only; and (iv) initial partitioning pool with KL refinement only. Figure 11 presents the results across all testcases from Table 4. The full Core-ChipletPart achieves an average cost reduction of 9.3% compared to the pool-only baseline, with improvements ranging from 6.4% to 13.4% across different designs. When examining the individual refinement strategies, FM-only achieves 6.4% average improvement while KL-only achieves 5.7% average improvement over the pool-only baseline. Notably, neither refinement strategy consistently dominates the other: FM-only outperforms KL-only on five out of ten testcases, while KL-only performs better on the remaining four testcases. This complementary behavior demonstrates that the two refinement algorithms optimize different aspects of the partitioning solution. FM excels at move-based local optimization while KL provides effective swap-based refinement for certain design topologies. Importantly, the full Core-ChipletPart consistently outperforms both individual refinement strategies across all testcases, achieving an additional 2-3% improvement over the better of FM-only or KL-only in each case. We believe that these results validate our integrated approach: the diverse initial partition pool provides high-quality starting points, while the application of FM and KL refinement generates further improvement that neither algorithm achieves independently.
@@ 6.2.4. Validation of heterogeneous technology-awareness
We evaluate benefits of incorporating heterogeneous technology awareness on all nine testcases (Table 4) and a technology list comprising 7 nm, 10 nm, and 14 nm nodes. Table 7 reports chiplet cost reductions achieved by enabling technology awareness, relative to the best solutions obtained under homogeneous technology integration for each node. Specifically, we first run ChipletPart using a single technology node per testcase (homogeneous integration) to obtain baseline partitioning solutions for each technology. We then run ChipletPart with technology awareness enabled (i.e., allowing the use of all three technology nodes) and compare the resulting solutions. Our results show that heterogeneous technology assignment can reduce system cost by up to 43%, with geometric mean cost reductions of 7%, 15%, and 15% when compared against homogeneous solutions at 7 nm, 10 nm, and 14 nm, respectively. [^24] To further validate the effectiveness of our GA, we implemented an enumeration-based framework that exhaustively explores all $\sum_{k=1}^{K_{\text{max}}}\binom{k+m-1}{m-1}$ possible assignments. Our experiments show that the GA achieves up to $7\times$ speedup with less than 1% degradation in chiplet cost compared to the enumeration-based framework. Note, that all solutions generated by ChipletPart are floorplan-feasible.
@@
Refer to caption Image includes 4 plots of the impact of hyperparameters on results.
@@
Refer to caption Image of a spider chart showing the sensitivity to various hyperparameters.
Figure 13. Sensitivity analysis of hyperparameters. Larger distances from the plot center indicate greater influence.
@@ 6.2.5. Hyperparameter exploration and sensitivity analysis
ChipletPart’s genetic algorithm relies on four key hyperparameters: (i) total population size $tot_{pop}$; (ii) maximum number of generations $\Psi$; (iii) mutation probability $p_{m}$; and (iv) crossover probability $p_{c}$ (cf. Algorithm 1). [^25] ChipletPart never returns a chiplet solution with more than 8 chiplets, so we set $K_{max}$ to 8. We validate the default values of these hyperparameters via an empirical study using testcases $WS_{1}$, $WS_{2}$, $TC_{1}$, and $TC_{2}$ from Table 4. We use the following technology nodes for this study: 7 nm, 10 nm, and 14 nm, and run ChipletPart on the four aforementioned testcases. Figure 12 presents the normalized chiplet costs (normalized to the cost of manual solutions) and the normalized runtime (represented as “Normalized Value” in Figure 12), normalized to a constant value of $10000$ seconds.2626261000010000 seconds is chosen for normalization to allow for comparison with about three hours of runtime. In our experiments, we systematically vary each hyperparameter while keeping the others fixed at their default values. For each configuration, we measure the resulting chiplet cost and runtime across the testcases. Our observations from Figure 12 indicates that the default hyperparameter settings achieve a balanced trade-off between partitioning quality and runtime efficiency.